# HMT: Hierarchical Memory Transformer for Long Context Language Processing

### v1

@@ 1 Introduction
  • HMT applies to any models with a short and limited context window to enable learning infinitely long context with enhanced effectiveness  [^1] In this paper, effectiveness means how well the language model generates meaningful texts. We measure it by perplexity, especially with context switching. We demonstrate that with HMT, OPT and OpenLlamaV2 can be 25.5% and 17.6% more effective in perplexity (PPL) on Wikitext-103 with multiple samples concatenated, and 11.4% and 9.48% on PG-19, a book dataset. We also evaluated HMT with Llama 2 7B regarding the PubMedQA dataset with multiple contexts. It was 9.81% more effective in long-answer contextual reasoning with a 1.0% higher short-answer prediction accuracy.
@@ 3.2 Memory Recall Mechanism
The overhead of the memory recall mechanism can be hidden if implemented on configurable hardware, as explained in Appendix F.
@@ 3.3 Hierarchical Memorization
Long-term Memory: Each generated memory embedding will be cached as the long-term memory. Based on the memory constraint of the hardware, HMT stores the most recent N memory embeddings. We denote these cached embeddings as $H^{mem}_{[n-N-1,n-1)}$. The cached embeddings will be utilized as the input of memory recall mechanism to generate the memorization token embedding $H^{S}_{n}$ for each segment as illustrated in the previous section.
@@ 4.1 Long-context Training
On the contrary, HMT will not suffer from gradient vanishing or explosion as BPTT unroll depth increases due to the memory recall mechanism. Table 2 reveals that HMT can improve its effectiveness continuously as the BPTT unroll depth increases during training. Therefore, HMT will be more effective when the BPTT unroll depth increases. A detailed gradient stability analysis is presented in Appendix G. Furthermore, we applied several techniques to optimize the GPU memory consumption to increase the maximum trainable BPTT unroll depth compared with RMT, described in Appendix H.
@@ 6.2 Comparison to Long Context Models
Furthermore, compared with other memory-augmented models, HMT is not only easy to use but also more effective. The Memorizing transformer (Wu et al., 2022) is a memory-augmented long-context model. This model requires training from scratch per downstream task before inference, which is not user-friendly for general usage. We compared the Memorizing Transformer trained on the Wikitext-103 dataset with HMT + OPT 350M, which has a similar parameter count, shown in table 5. With only 700 training steps, HMT outperforms the Memorizing Transformer trained with 10k steps by 2x in perplexity, indicating that HMT is more efficient and effective to use.
@@ 6.3 Comparison to RMT
For recurrent sequence models, HMT can further increase the effectiveness of RWKV while RMT will degrade the performance for both datasets, as demonstrated in Figure 5. Since RWKV has already compressed past tokens and passed hidden states along the sequence, applying RMT to RWKV is simply re-weighting past information compressed in states periodically. This was originally done by the time-mixing module of RWKV. Therefore, the advantage of memory augmentation is limited. Due to the gradient vanishing issue mentioned in Appendix G, the model is harder to train with RMT, leading to inferior performance. However, we believe that the memory recall mechanism in HMT helps RWKV to select previous hidden states with the most relevance, boosting the effectiveness.
@@ 6.4 Ablation Study
Impact of summarizing partial segment in memory recall. To overlap or reduce the inference time of the previous segment with the representation extraction of the next segment, it is necessary to prefetch only the first $l$ tokens in the segment for summarization as described in Appendix F. In the experiment, we select half of the segment for representation extraction. We examine the model that extracts the whole segment and compare the effectiveness, depicted by Figure 9. The impact is negligible. We hypothesize that the start of a segment contains enough information about the overall topic for memory recall, which is intuitive as humans can capture the concepts by browsing keywords or segments instead of reading the whole paragraphs.
@@ Appendix A Other Related Works
Appendix B Comparison to Unlimiformer
@@
Appendix C Ablation Study of HMT Memory Recall Behavior
@@
Appendix D Relationships Between Effectiveness and Size of Sensory Memory
@@
Appendix E Dataset Split for PubMedQA
@@
Appendix F Hardware Acceleration Consideration
@@
Appendix G Gradient Stability in HMT and RMT
@@
Appendix H Distributed Training with Memory Consumption Optimization

### v2

@@ 1 Introduction
  • HMT applies to any models with a short and limited context window to enable learning infinitely long context with enhanced effectiveness  [^1] In this paper, effectiveness means how well the language model generates meaningful texts. It is measured in perplexity., especially with context switching. We demonstrate that with HMT, OPT and OpenLlamaV2 can be 25.5% and 17.6% more effective in perplexity (PPL) on Wikitext-103 with multiple samples concatenated, and 11.4% and 9.48% on PG-19, a book dataset. We also evaluated HMT with Llama 2 7B regarding the PubMedQA dataset with multiple contexts. It was 9.81% more effective in long-answer contextual reasoning with a 1.0% higher short-answer prediction accuracy.
@@ 3.2 Memory Recall Mechanism
The overhead of the memory recall mechanism can be hidden if implemented on configurable hardware, as explained in Appendix G.
@@ 3.3 Hierarchical Memorization
Long-term Memory: Each generated memory embedding will be cached as the long-term memory. Based on the memory constraint of the hardware, HMT stores the most recent N memory embeddings. The cached embeddings will be utilized as the input of memory recall mechanism to generate the memorization token embedding $H^{S}_{n}$ for each segment as illustrated in the previous section.
@@ 4.1 Long-context Training
On the contrary, HMT will not suffer from gradient vanishing or explosion as BPTT unroll depth increases due to the memory recall mechanism. Table 2 reveals that HMT can improve its effectiveness continuously as the BPTT unroll depth increases during training. Therefore, HMT will be more effective when the BPTT unroll depth increases. A detailed gradient stability analysis is presented in Appendix H. Furthermore, we applied several techniques to optimize the GPU memory consumption to increase the maximum trainable BPTT unroll depth compared with RMT, described in Appendix I.
@@ 6.2 Comparison to Long Context Models
Furthermore, compared with other memory-augmented models, HMT is not only easy to use but also more effective. The Memorizing transformer (Wu et al., 2022) is a memory-augmented long-context model. This model requires training from scratch per downstream task before inference, which is not user-friendly for general usage. We compared the Memorizing Transformer trained on the Wikitext-103 dataset with HMT + OPT 350M, which has a similar parameter count, shown in table 5. With only 700 training steps, HMT outperforms the Memorizing Transformer trained with 10k steps by 2x in perplexity, indicating that HMT is more efficient and effective to use. We also compared HMT with another memory-augmented model, LongMem (Wang et al., 2024), shown in Appendix B.
@@ 6.3 Comparison to RMT
For recurrent sequence models, HMT can further increase the effectiveness of RWKV while RMT will degrade the performance for both datasets, as demonstrated in Figure 5. Since RWKV has already compressed past tokens and passed hidden states along the sequence, applying RMT to RWKV is simply re-weighting past information compressed in states periodically. This was originally done by the time-mixing module of RWKV. Therefore, the advantage of memory augmentation is limited. Due to the gradient vanishing issue mentioned in Appendix H, the model is harder to train with RMT, leading to inferior performance. However, we believe that the memory recall mechanism in HMT helps RWKV to select previous hidden states with the most relevance, boosting the effectiveness.
@@ 6.4 Ablation Study
Impact of summarizing partial segment in memory recall. To overlap or reduce the inference time of the previous segment with the representation extraction of the next segment, it is necessary to prefetch only the first $l$ tokens in the segment for summarization as described in Appendix G. In the experiment, we select half of the segment for representation extraction. We examine the model that extracts the whole segment and compare the effectiveness, depicted by Figure 9. The impact is negligible. We hypothesize that the start of a segment contains enough information about the overall topic for memory recall, which is intuitive as humans can capture the concepts by browsing keywords or segments instead of reading the whole paragraphs.
@@ Appendix A Other Related Works
Appendix B Comparison to LongMem
Unlike HMT, LongMem (Wang et al., 2024) operates on the key and value caches of each layer of the model and requires caching long caches to capture distant context. To compare with LongMem, we pick the Qwen1.5-0.5B (Bai et al., 2023) model as the backbone model and train HMT by 700 steps with 4 segments over the ArXiv subset of the Pile dataset (Gao et al., 2020), which is the same setting as LongMem. Due to the large storage consumption of the training subset of the Pile dataset, we only extract the ArXiv subset in the validation and test split. HMT is trained on the validation set and tested on the test set. Table 6 illustrates that HMT + Qwen1.5-0.5B achieves lower PPL, with a smaller parameter size and in-memory length (number of memory embeddings, which is the number of key and value embeddings cached for LongMem). This indicates that HMT is memory efficient.
Table 6: Effectiveness of LongMem and HMT + Qwen1.5-0.5B models over ArXiv subset of the Pile dataset. With HMT, the Qwen1.5-0.5B model can obtain better effectiveness with fewer parameters and shorter memory, after 700 steps of update. The result for LongMem comes from the original paper.
Model # Params Segment Length In-memory Length Test PPL (ArXiv)
LongMem 558M 1k 65k 10.08
HMT + Qwen1.5-0.5B 463M 1k 300 9.06
Appendix C Comparison to Unlimiformer
@@
Appendix D Ablation Study of HMT Memory Recall Behavior
@@
Appendix E Relationships Between Effectiveness and Size of Sensory Memory
@@
Appendix F Dataset Split for PubMedQA
@@
Appendix G Hardware Acceleration Consideration
@@
Appendix H Gradient Stability in HMT and RMT
@@
Appendix I Distributed Training with Memory Consumption Optimization